Grade 12th PassGeneral Physics

a galvanometer has resistance of 10 ohm and gives a full scale deflection when a current of 0.001 amp flow in it. find out the value of the shunt resistance to convert it into an ammeter of range 10 amperes

Saiba Muhammed
7 Years agoGrade 12th Pass

1 Answer

Khimraj
7 Years ago
Ig = IS/(G + S) 
Ig*G + Ig*S = IS
S = Ig*G/(I -Ig)
S = 0.001*10/(10- 0.001) = 0.001 ohm
Hope it clears...........................